OpenStudy (anonymous):

log base 7 (x+2)-log base 7(x)=2 help me solve this

OpenStudy (anonymous):

I need helpppp

OpenStudy (anonymous):

when you subtract logs the y become a division problem. so\[\log_{7}(X+2)-\log_{7}x=\log_{7}((x+2)/x) \] now you just have to chnge the bases and solve

OpenStudy (anonymous):

\[\log_{7}(x+2/x)=\log(x+2/x)/\log7 \]

OpenStudy (anonymous):

Are you there?